OpenStudy (anonymous):

How many moles of chlorine gas cl2 are contained in a 10.3 l tank at 21.2

OpenStudy (australopithecus):

Just use the ideal gas equation look on youtube or go on wikipedia

OpenStudy (anonymous):

The units in the gas constant tell you what units you need to use, because everything, except the number of moles, must cancel out. Since pressure is given in torr (mm Hg), it must be converted to atm. 633 torr/760 torr/atm = 0.833 atm. Furthermore, you can't use degrees celcius, which must be converted to Kelvin (add 273.15 degrees - we'll round to 273.2 degrees to keep it simple): T = 273.2 + 21.2 = 294 K (three significant figures). So the number of moles n = PV/RT = 0.833 atm x 10.3 L /0.0821 L atm/mol K x 294 K = 0.355 moles